# Iwahori Matsumoto presentation for modules of Iwahori fixed functions on symmetric spaces

### v2

@@ 4 The Key Computation
Let $X=G/H$ be a $p$-adic symmetric space, recall that our goal is to describe the action of the Iwahori Hecke algebra, $\mathcal{H}(G,I)$ on $S(X)^{I}$. By Theorem 2.32 it is enough to describe the action of $T_{s}$ for $s\in\tilde{\Delta}$ and the action of $T_{o}$ for $o\in\Omega$.
We begin with describing the action of $T_{o}$ for $o\in\Omega$.
@@
Let $o\in\Omega$ and let $x\in I\backslash G/H$. Let $T_{o}\in\mathcal{H}(G,I)$ be the characteristic function of $o$ and let $1_{x}\in S(X)^{I}$ be the characteristic function of $x$. Then $T_{o}1_{x}=1_{o\times x}$
@@
Let $m$ be a left Haar measure on $G$ normalized such that $m(I)=1$.
[…]
If $T_{o}1_{x}(y)\neq 0$ then there is $g\in IoI$ such that $g^{-1}y\in IxH$, therefore $y\in IoIxH$. $\Omega$ normalizes $I$ so we must have $y\in IoxH$. Thus, $T_{o}1_{x}$ is proportional to $1_{o\times x}$. If $y\in IoxH$ then $T_{o}1_{x}(y)=m(IoI)=1$ so we are done. ∎
Let $x\in I\backslash G/H$ and let $s\in\tilde{\Delta}$ be a simple reflection.
@@
In remaining part of this Section we will translate the problem of computing $T_{s}1_{x}$ to a combinatorial problem about $H$ orbits of chambers in $\mathcal{B}_{G}$, the building of $G$.
@@
Let $\mathcal{C}$ be the chamber fixed by $I$ and let $\mathcal{A}$ be an apartment that contains $\mathcal{C}$, let $s\in\tilde{\Delta}$ be an element that induces a reflection over a face of $\mathcal{C}$.
The set of chambers $Is\mathcal{C}$ is the set of all chambers that contain $\mathcal{C}\cap s\mathcal{C}$ which are different from $\mathcal{C}$ and its size is $q=\#k$.
@@
Now let $\mathcal{C}_{1}$ be a chamber that contains $f$ and is different from $\mathcal{C}$ and $s\mathcal{C}$. By theorem 2.12 there is an apartment $\mathcal{A}^{\prime}$ that contains $\mathcal{C}$ and $\mathcal{C}_{1}$, also there is $g\in G$ such that $g\mathcal{A}=\mathcal{A}^{\prime}$ and such that $g$ fixes $\mathcal{A}\cap\mathcal{A}^{\prime}$ pointwise. $\mathcal{C}\subset\mathcal{A}\cap\mathcal{A}^{\prime}$ so $g$ fixes $\mathcal{C}$ and it sends $s\mathcal{C}$ to $\mathcal{C}^{\prime}$. We found $g\in I$ such that $gs\mathcal{C}=\mathcal{C}_{1}$ which proves $\mathcal{C}_{1}\in Is\mathcal{C}$.
It is well known that chambers that contain a given face $f$ corresponds bijectivly to $\mathbb{P}^{1}(k)$ (for example, this follows from Propositions 2.25 and 2.29). Excluding $\mathcal{C}$ there are $q$ chambers that contain $f$.
Let $\mathcal{C}$ be the chamber fixed pointwise by $I$, let $g\in G$ be a representative of the orbit $x$.
There is $o\in\Omega$ such that $(g\mathcal{C},o)$ is a $\Omega$ colored chamber whose $H$ orbit is $x$, $(gs\mathcal{C},o)$ is a $\Omega$ colored chamber whose orbit is $s\times x$, denote $f=g\mathcal{C}\cap gs\mathcal{C}$.
@@
Let $s\in\tilde{\Delta}$, $x\in I\backslash G/H$, $g$ a representative of $x$, $f=g\mathcal{C}\cap gs\mathcal{C}$ and $\mathcal{C}_{1},...,\mathcal{C}_{q+1}$ are the chambers that contain $f$. Denote $o=\omega(g)\in\Omega$
@@
For each $11\leq i\leq q+$ 1 choose $g_{1},...,g_{q+1}\in G$ such that $g_{i}\mathcal{C}_{0}=\mathcal{C}_{i}$ and $\omega(g_{i})=\omega(g)$. Denote $x_{i}=g_{i}H$.
We want to compute $T_{s}1_{x}(x_{i})$. Let $m$ be a left Haar measure on $G$ normalized such that $m(I)=1.$ A direct calculation gives:
[…]
On other hand,
$m(\{k\in IsI|k^{-1}g_{i}\in IgH\})=m(IsI\cap g_{i}Hg^{-1}I)=m(g^{-1}_{i}IsI\cap Hg^{-1}I).$
To calculate the intersection, we use lemma 4.2 to write $IsIg_{i}=\cup_{k\neq i}Ig_{k}$
This clearly implies that
$g^{-1}_{i}IsI=\cup_{k\neq i}g^{-1}_{k}I$
@@
$T_{s}1_{x}(x_{i})=\sum_{k\neq i}m(g^{-1}_{k}I\cap Hg^{-1}I)$
We observe that by the our choice of $g_{1},...,g_{q+1},$ for every $k$ we either have $g_{k}^{-1}I\subset Hg^{-1}I$ or $g_{k}^{-1}I\cap Hg^{-1}I=\varnothing.$
@@
$\beta_{g\mathcal{C},g_{i}\mathcal{C}}:=\sum_{k\neq i}\delta_{Hg_{k}I,HgI}$
@@
$T_{s}1_{x}(x_{i})=\sum_{k\neq i}\delta_{Hg_{k}I,HgI}$
motivate us to consider the quantity $T_{s}1_{x}(x_{i})+1_{x}(x_{i}).$ which is independent of $i$ and equals to $\gamma_{f,g}$.
Thus we obtain $(T_{s}+1)1_{x}(x_{i})=\gamma_{f,g}$.
This computation also shows that $(T_{s}+1)1_{x}$ is supported on $\{x_{1},...,x_{q+1}\}$ so we must have $(T_{s}+1)1_{x}=\gamma_{f,g}D_{f,o}$
We see that in order to compute $T_{s}1_{x}$, where $x\in I\backslash G/H$ and $s\in\tilde{\Delta}$ is a simple reflection we need to answer the following combinatorial question: given $s$ as above, a chamber $\mathcal{C}$ and $o\in\Omega$ such that $(\mathcal{C},o)$ corresponding to $x$, determine the sizes of the $H$ orbits of chambers that contain $f=\mathcal{C}\cap s\mathcal{C}$ with the color $o\in\Omega$. We will address this in the following sections.
@@ 5 Iwahori-Matsumoto type formulas
In this section we prove a result that generalizes the classical Iwahori-Matsumoto description given in theorem 2.31, thus giving a partial solution to the problem of describing the action of $\mathcal{H}(G,I)$ on $S(X)^{I}$. This allows us to show that $S(X)^{I}$ is finitely generated over $\mathcal{H}(G,I)$.
In the description of Iwahori-Matsumoto, a key role is played by a length function on the affine Weyl group. For our description, we will use a length function $\ell_{\sigma}$ on the set $I\backslash G/H.$
@@ 5.1 The length function
Our length function will be defined first on the chambers of the building. We begin by introducing this length function, $\ell_{\sigma}$.
Recall from definition 2.17, for any two chambers $\mathcal{C}_{1},\mathcal{C}_{2}\subset\mathcal{B}_{G}$, their distance $d(\mathcal{C}_{1},\mathcal{C}_{2})$ is defined to be the length of a minimal gallery between them.
@@
$l_{\sigma}(\mathcal{C}):=d(\mathcal{C},\sigma(\mathcal{C}))$
@@
Let $h\in H$ and $\mathcal{C}$ a chamber, then $l_{\sigma}(h\mathcal{C})=d(h\mathcal{C},\sigma(h\mathcal{C}))=d(h\mathcal{C},h\sigma(\mathcal{C}))=d(\mathcal{C},\sigma(\mathcal{C}))$. ∎
We use the notation $\ell_{\sigma}$ also for the induced length function on the $H$ orbits on the chambers of the building.
Under the identification between $I\backslash G/H$ with $H$ orbits of $\Omega$ colored chambers in the building, $l_{\sigma}$ induces a length function on $I\backslash G/H$ which we also denote by $l_{\sigma}$.
For $s\in\tilde{\Delta}$ a simple reflection let $T_{s}\in\mathcal{H}(G,I)$ be the corresponding element of the Hecke algebra.
@@
For $x\in I\backslash G/H$ let $1_{x}\in S(X)^{I}$ be the characteristic function of the orbit $x$. Let $s\in\tilde{\Delta}$ be a simple reflection.
@@
  • If $l_{\sigma}(s\times x)>l_{\sigma}(x)$ then $T_{s}1_{x}=1_{s\times x}$.
  • If $l_{\sigma}(s\times x)<l_{\sigma}(x)$ then $T_{s}1_{x}=(q-1)1_{x}+q1_{s\times x}$.
@@
In the group case it is always true that $l_{\sigma}(s\times x)\neq l_{\sigma}(x)$, thus the theorem indeed generalizes the classical description of Iwahori-Matsumoto. However, for a general symmetric space it might happen that $l_{\sigma}(s\times x)=l_{\sigma}(x)$. The general case will be addressed in theorem 7.12.
@@
Let $x\in I\backslash G/H,s\in\tilde{\Delta}$ a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$. Choose $g\in G$ and choose a $\sigma$-stable apartment $\mathcal{A}$ such that $g$ is representative of the orbit $x$ and $g\mathcal{C},gs\mathcal{C}\subset\mathcal{A}$ (see Remark 3.4) and let $f=g\mathcal{C}\cap gs\mathcal{C}$.
@@
  • $l_{\sigma}(s\times x)>l_{\sigma}(x)$ if and only if $g\mathcal{C}\subset\Pi$.
  • $l_{\sigma}(s\times x)<l_{\sigma}(x)$ if and only if $gs\mathcal{C}\subset\Pi$.
@@
Let us look at all the $\sigma$ stable apartments that contain $f$, there are 3 cases:
@@
    Neither case 1 nor case 2 happen.
We will show first that in case 3 we necessarily have $l_{\sigma}(s\times x)=l_{\sigma}(x)$ and thus nothing to prove.
In case 3 the intersection $\Pi$, of all $\sigma$ stable apartments that contain $f$, is a convex set in $\mathcal{A}$ (by theorem 2.16) that contains $f$ but doesn’t contain any chamber that contains $f$ so it is contained in $P$, the affine hyper-plane spanned by $f$.
$P$ is $\sigma$ stable as it is spanned by the intersection of $\sigma$ stable apartments.
@@
Let $(\mathcal{C}_{i})_{i=1}^{m}$ be a minimal gallery that starts with $\mathcal{C}_{1}=g\mathcal{C}$ and ends with $\mathcal{C}_{m}=\sigma(g\mathcal{C}).$ Observe that by theorem 2.18 these chambers are all contained in the apartment $\mathcal{A}.$ Applying the reflection $r_{P}$ we obtain a minimal gallery $(r_{P}(\mathcal{C}_{i}))_{i=1}^{m}$ that starts with $r_{P}(g\mathcal{C}_{1})=gs\mathcal{C}$ and ends with $r_{P}(\sigma(g\mathcal{C}))=\sigma(r_{P}(g\mathcal{C}))=\sigma(gs\mathcal{C})$ which means that $l_{\sigma}(s\times x)=l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})=l_{\sigma}(x)$.
This shows that in case 3 we have $l_{\sigma}(s\times x)=l_{\sigma}(x)$. For the other cases it is enough to show that if $l_{\sigma}(gs\mathcal{C})>l_{\sigma}(g\mathcal{C})$ then case 1 happens, that is all the $\sigma$ stable apartments that contain $f$ also contain $g\mathcal{C}$.
If $l_{\sigma}(gs\mathcal{C})>l_{\sigma}(g\mathcal{C})$ then $l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})+2$ as $l_{\sigma}$ is constant modulo $2$ on each apartment. This means that we can choose a minimal gallery of chambers from $g\mathcal{C}$ to $\sigma(g\mathcal{C})$ and then add $gs\mathcal{C}$ at the start and $\sigma(gs\mathcal{C})$ at the end and get a minimal gallery of chambers from $gs\mathcal{C}$ to $\sigma(gs\mathcal{C})$. By this and theorem 2.18 any $\sigma$ stable apartment that contains $gs\mathcal{C}$ has to contain $g\mathcal{C}$. Therefore either case 1 or case 3 happen and we know that it isn’t case 3.
@@
Let $\mathcal{C}$ be the chamber fixed by $I$, let $g\in G$ a representative of the orbit $x$ and let $\mathcal{A}$ be a $\sigma$ stable apartment that contains $g\mathcal{C}$ and $gs\mathcal{C}$.
If $l_{\sigma}(gs\mathcal{C})>l_{\sigma}(g\mathcal{C})$ then by lemma 5.5 all $\sigma$ stable apartments that contain $f=g\mathcal{C}\cap gs\mathcal{C}$ must contain $g\mathcal{C}$.
@@
This means that all chambers that contain $f$ and are different from $g\mathcal{C}$ are in the $H$ orbit of $gs\mathcal{C}$. $g\mathcal{C}$ it self is not in this $H$ orbit. By proposition 4.3 this yields $(T_{s}+1)1_{x}=1_{s\times x}+1_{x}$, or $T_{s}1_{x}=1_{s\times x}$
@@
Let $x\in I\backslash G/H$ be like in the previous theorem, let $w\in W_{aff}(T)$ and let $l:W_{aff}\rightarrow\mathbb{N}$ be the usual length function on $W_{aff}$ (see Section 2.1 for a definition).
If $l_{\sigma}(w\times x)=2\cdot l(w)+l_{\sigma}(x)$ then
$T_{w}1_{x}=1_{w\times x}$
@@
Write $w=s_{1},...,s_{l(w)}o$ as a product of simple reflections $s_{i}\in\tilde{\Delta}$ and $o\in\Omega$.
By theorem 2.31 $T_{w}=T_{s_{1}}...T_{s_{l(w)}}T_{o}$, for any $1\leq j\leq l(w)$ we have $l_{\sigma}(s_{j}\times...\times s_{l(w)}\times o\times x)\leq l_{\sigma}(x)+2(l(w)-j+1)$ and we must have an equality otherwise we wouldn’t have the equality in the assumption as acting by each $s_{i}$ can increase the length by at most $2$.
By theorem 5.3, proposition 4.1 and induction we get that $T_{s_{j}}...T_{s_{l(w)}}T_{o}1_{x}=1_{s_{j}\times...\times s_{l(w)}\times o\times x}.$ For $j=1$ we obtain the required result. ∎
This result is enough to prove that $S(X)^{I}$ is finitely generated over ${\cal H}(G,I).$
@@
Let $T$ be a $\sigma$ stable maximal split torus of $G$ and let $\mathcal{A}$ be the apartment that corresponds to $T$. $T$ acts on $\mathcal{A}$ by translations. Let $t\in T$ such that its action on $\mathcal{A}$ commutes with $\sigma$, then there is $t^{\prime}\in T\cap H$ whose action on $\mathcal{A}$ is the same as $t$.
@@
Let $\pi$ be a uniformaizer of $\mathcal{O}$ and let $\Lambda(T)=\{\chi^{\vee}(\pi)|\chi^{\vee}\in X^{*}(T)\}$. $T$ is a direct product of $\Lambda(T)$ and $T^{0}$ (See 3.6 in [DS11]). Write $t=t_{1}t_{2}$ with $t_{1}\in\Lambda(T)$ and $t_{2}\in T^{0}$.
Choose a chamber $\mathcal{C}\subset\mathcal{A}$, we have $\sigma(t)\sigma(\mathcal{C})=\sigma(t\mathcal{C})=t\sigma(\mathcal{C})$, so $t^{-1}\sigma(t)$ acts as a translation that fixes $\sigma(\mathcal{C})$ and thus it acts trivially, meaning $t^{-1}\sigma(t)\in T^{0}$. Thus $t_{1}^{-1}\sigma(t_{1})\in T^{0}$. $\Lambda(T)$ is $\sigma$ stable so we must have $t^{-1}_{1}\sigma(t_{1})\in T^{0}\cap\Lambda(T)=1$ or $t_{1}\in H$. $t$ and $t_{1}$ act the same way on $\mathcal{A}$ so we are done. ∎
@@
Let $\mathcal{A}$ be a $\sigma$ stable apartment, let $d\in\mathbb{N}$, the number of $H$ orbits of chambers $\mathcal{C}$ in $\mathcal{A}$ with $l_{\sigma}(\mathcal{C})=d$ is finite.

### v3

@@ 4 The Key Computation
[…]
We begin by describing the action of $T_{o}$ for $o\in\Omega$.
@@
Let $o\in\Omega$ and let $x\in H\backslash G/I$. Let $T_{o}\in\mathcal{H}(G,I)$ be the characteristic function of $o$ and let $1_{x}\in S(X)^{I}$ be the characteristic function of $x$. Then $1_{x}T_{o}=1_{x\times o}$.
@@
Let $m$ be a Haar measure on $G$ normalized so that $m(I)=1$.
$1_{x}T_{o}(y)=\int_{g\in G}1_{x}(yg^{-1})T_{o}(g)dm(g)$
If $1_{x}T_{o}(y)\neq 0$, then there is $g\in IoI$ such that $yg^{-1}\in HxI$, therefore $y\in HxIoI$. The group $\Omega$ normalizes $I$ so we obtain $y\in HxoI$. Thus, $1_{x}T_{o}$ is proportional to $1_{x\times o}$. If $y\in HxoI$ then $1_{x}T_{o}(y)=m(IoI)=1$, this completes the proof. ∎
Let $x\in H\backslash G/I$ and let $s\in\tilde{\Delta}$ be a simple reflection.
@@
In the remainder of this section, we translate the problem of computing $1_{x}T_{s}$ to a combinatorial problem about $H$ orbits of chambers in $\mathcal{B}_{G}$, the building of $G$.
@@
Let $\mathcal{C}$ be the chamber fixed by $I$ and let $\mathcal{A}$ be an apartment that contains $\mathcal{C}$. Let $s\in\tilde{\Delta}$ be an element that induces a reflection over a facet of $\mathcal{C}$.
The set of chambers $Is\mathcal{C}$ is the set of all chambers that contain $\mathcal{C}\cap s\mathcal{C}$ and are different from $\mathcal{C}$ and its size is $q$, the size of the residue field of $F$.
@@
Now, let $\mathcal{C}_{1}$ be a chamber that contains $f$ and is different from both $\mathcal{C}$ and $s\mathcal{C}$. By Theorem 2.12 there is an apartment $\mathcal{A}^{\prime}$ that contains both $\mathcal{C}$ and $\mathcal{C}_{1}$. Also there is $g\in G$ such that $g\mathcal{A}=\mathcal{A}^{\prime}$ and such that $g$ fixes $\mathcal{A}\cap\mathcal{A}^{\prime}$ pointwise. We have $\mathcal{C}\subset\mathcal{A}\cap\mathcal{A}^{\prime}$ so $g$ fixes $\mathcal{C}$ pointwise and it sends $s\mathcal{C}$ to $\mathcal{C}^{\prime}$. We found $g\in I$ such that $gs\mathcal{C}=\mathcal{C}_{1}$ which proves $\mathcal{C}_{1}\in Is\mathcal{C}$.
It is well known that the set of chambers that contain a given facet $f$ corresponds bijectively to $\mathbb{P}^{1}(k)$ (for example, this follows from Propositions 2.26 and 2.30). Excluding $\mathcal{C}$ there are $q$ chambers that contain $f$.
Let $\mathcal{C}$ be the chamber fixed pointwise by $I$ and let $g\in G$ be a representative of the orbit $x$ such that $g\mathcal{A}_{0}$ is a $\sigma$ stable apartment.
There exists $o\in\Omega$ such that $(g\mathcal{C},o)$ is a $\Omega$ colored chamber whose $H$ orbit is $x$, and $(gs\mathcal{C},o)$ is a $\Omega$ colored chamber whose $H$ orbit is $x\times s$, denote $f=g\mathcal{C}\cap gs\mathcal{C}$.
@@
Let $s\in\tilde{\Delta}$, $x\in H\backslash G/I$, $g$ a representative of $x$ as above. Let $f=g\mathcal{C}\cap gs\mathcal{C}$ and let $\mathcal{C}_{1},...,\mathcal{C}_{q+1}$ be the chambers that contain $f$. Denote $o=\omega(g)\in\Omega$.
@@
First, observe that by the definition of convolution, $1_{x}(T_{s}+1)$ is supported on $HxIsI$. And that all the $H$ orbits of chambers in $HxIs\mathcal{C}$ can be represented by a chamber that contains $f$.
For each $1\leq i\leq q+1$ choose $g_{1},...,g_{q+1}\in G$ such that $g_{i}\mathcal{C}_{0}=\mathcal{C}_{i}$ and $\omega(g_{i})=\omega(g)$. Denote $x_{i}=Hg_{i}$.
We now compute $1_{x}T_{s}(x_{i})$. Let $m$ be a Haar measure on $G$ normalized such that $m(I)=1.$ A direct calculation gives:
$1_{x}T_{s}(x_{i})=\int_{g^{\prime}\in G}1_{x}(x_{i}g^{\prime-1})T_{s}(g^{\prime})dm(g^{\prime})=m(\{k\in IsI|g_{i}k^{-1}\in HgI\}).$
On the other hand,
$m(\{k\in IsI|g_{i}k^{-1}\in HgI\})=m(IsI\cap Ig^{-1}Hg_{i})=m(IsIg^{-1}_{i}\cap Ig^{-1}H).$
To compute the intersection, we use Lemma 4.2 to write $g_{i}IsI=\cup_{k\neq i}g_{k}I$
Taking inverses, we obtain
$IsIg^{-1}_{i}=\cup_{k\neq i}Ig^{-1}_{k}$
@@
$1_{x}T_{s}(x_{i})=\sum_{k\neq i}m(Ig^{-1}_{k}\cap Ig^{-1}H)$
We observe that, by the our choice of $g_{1},...,g_{q+1},$ for every $k$ we either have $Ig_{k}^{-1}\subset Ig^{-1}H$ or $Ig_{k}^{-1}\cap Ig^{-1}H=\varnothing.$
@@
$\beta_{g\mathcal{C},g_{i}\mathcal{C}}=\sum_{k\neq i}\delta_{Hg_{k}I,HgI}$
@@
$1_{x}T_{s}(x_{i})=\sum_{k\neq i}\delta_{Hg_{k}I,HgI}$
motivates us to consider the quantity $1_{x}T_{s}(x_{i})+1_{x}(x_{i}).$ which is independent of $i$ and equals $\gamma_{f,g}$.
Thus we obtain $1_{x}(T_{s}+1)(x_{i})=\gamma_{f,g}$.
This computation also shows that $1_{x}(T_{s}+1)$ is supported on $\{x_{1},...,x_{q+1}\}$ so we obtain $1_{x}(T_{s}+1)=\gamma_{f,g}D_{f,o}$
Remark 4.4.
Notice that the value $\gamma_{f,g}$ depends on the chamber $g\mathcal{C}$. For different chambers among $\mathcal{C}_{1},...,\mathcal{C}_{q+1}$ the value may vary.
We see that to compute $1_{x}T_{s}$, for $x\in H\backslash G/I$ and $s\in\tilde{\Delta}$ a simple reflection, we need to answer the following combinatorial question. Given $s$ as above, a chamber $\mathcal{C}$ and $o\in\Omega$ such that $(\mathcal{C},o)$ corresponds to $x$, determine the sizes of the $H$ orbits of chambers that contain $f=\mathcal{C}\cap s\mathcal{C}$ with the color $o\in\Omega$. We address this in the following sections.
@@ 5 Iwahori-Matsumoto type formulas
In this section, we prove a result that generalizes the classical Iwahori-Matsumoto description given in Theorem 2.32. Thus, giving a partial solution to the problem of describing the action of $\mathcal{H}(G,I)$ on $S(X)^{I}$. This allows us to show that $S(X)^{I}$ is finitely generated over $\mathcal{H}(G,I)$.
[…]
@@ 5.1 The length function
Our length function will be defined first on the chambers of the building. We begin by introducing this length function, $l_{\sigma}$.
For any two chambers $\mathcal{C}_{1},\mathcal{C}_{2}\subset\mathcal{B}_{G}$, their distance $d(\mathcal{C}_{1},\mathcal{C}_{2})$ is defined to be the length of a minimal gallery between them.
@@
$l_{\sigma}(\mathcal{C})=d(\mathcal{C},\sigma(\mathcal{C}))$
@@
Let $h\in H$ and $\mathcal{C}$ be a chamber, then $l_{\sigma}(h\mathcal{C})=d(h\mathcal{C},\sigma(h\mathcal{C}))=d(h\mathcal{C},h\sigma(\mathcal{C}))=d(\mathcal{C},\sigma(\mathcal{C}))$. ∎
We use the notation $l_{\sigma}$ also for the induced length function on the $H$ orbits of chambers and on the $H$ orbits of $\Omega$ colored chambers.
[…]
For $s\in\tilde{\Delta}$ a simple reflection, let $T_{s}\in\mathcal{H}(G,I)$ be the corresponding element of the Hecke algebra.
@@
For $x\in H\backslash G/I$ let $1_{x}\in S(X)^{I}$ be the characteristic function of the orbit $x$. Let $s\in\tilde{\Delta}$ be a simple reflection.
@@
  1. 1.
    If $l_{\sigma}(x\times s)>l_{\sigma}(x)$ then $1_{x}T_{s}=1_{x\times s}$.
  2. 2.
    If $l_{\sigma}(x\times s)<l_{\sigma}(x)$ then $1_{x}T_{s}=(q-1)1_{x}+q1_{x\times s}$.
@@
In the group case it is always true that $l_{\sigma}(x\times s)\neq l_{\sigma}(x)$, thus the theorem indeed generalizes the classical description of Iwahori and Matsumoto. However, for a general symmetric space, it may happen that $l_{\sigma}(x\times s)=l_{\sigma}(x)$. The general case will be addressed in Theorem 7.13.
@@
Let $x\in H\backslash G/I$ and let $s\in\tilde{\Delta}$ be a simple reflection. Let $\mathcal{C}$ be the chamber fixed by $I$. Choose $g\in G$ and choose a $\sigma$-stable apartment $\mathcal{A}$ such that $g$ is a representative of the orbit $x$ and $g\mathcal{C},gs\mathcal{C}\subset\mathcal{A}$ (see Remark 3.7). Let $f=g\mathcal{C}\cap gs\mathcal{C}$.
@@
  1. 1.
    $l_{\sigma}(x\times s)>l_{\sigma}(x)$ if and only if $g\mathcal{C}\subset\Pi$.
  2. 2.
    $l_{\sigma}(x\times xs)<l_{\sigma}(x)$ if and only if $gs\mathcal{C}\subset\Pi$.
@@
Consider all the $\sigma$ stable apartments that contain $f$, there are three cases:
@@
    Neither case 1 nor case 2 happens.
We will begin by showing that in case 3 we necessarily have $l_{\sigma}(x\times s)=l_{\sigma}(x)$ and thus there is nothing to prove.
The intersection of all $\sigma$ stable apartments that contain $f$ is a convex set in $\mathcal{A}$ (by Theorem 2.17) that contains $f$. In case 3, $\Pi$ does not contain any chamber in $\mathcal{A}$ that contains $f$ so $\Pi$ is contained in $P$, the affine hyper-plane spanned by $f$.
$P$ is $\sigma$ stable as it is spanned by $\Pi$, the intersection of $\sigma$ stable apartments.
@@
Let $(\mathcal{C}_{i})_{i=1}^{m}$ be a minimal gallery that starts at $\mathcal{C}_{1}=g\mathcal{C}$ and ends at $\mathcal{C}_{m}=\sigma(g\mathcal{C}).$ Observe that by Theorem 2.19 these chambers are all contained in the apartment $\mathcal{A}.$ Applying the reflection $r_{P}$ we obtain a minimal gallery $(r_{P}(\mathcal{C}_{i}))_{i=1}^{m}$ that starts with $r_{P}(g\mathcal{C}_{1})=gs\mathcal{C}$ and ends with $r_{P}(\sigma(g\mathcal{C}))=\sigma(r_{P}(g\mathcal{C}))=\sigma(gs\mathcal{C})$ which means that $l_{\sigma}(x\times s)=l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})=l_{\sigma}(x)$.
This shows that in case 3 we have $l_{\sigma}(x\times s)=l_{\sigma}(x)$.
To handle the other cases, it suffices to show that if $l_{\sigma}(gs\mathcal{C})>l_{\sigma}(g\mathcal{C})$ then case 1 occurs. That is, all $\sigma$ stable apartments that contain $f$ also contain $g\mathcal{C}$.
If $l_{\sigma}(gs\mathcal{C})>l_{\sigma}(g\mathcal{C})$ then $l_{\sigma}(gs\mathcal{C})=l_{\sigma}(g\mathcal{C})+2$ as $l_{\sigma}$ is constant modulo $2$ on each apartment. This means that we may choose a minimal gallery of chambers from $g\mathcal{C}$ to $\sigma(g\mathcal{C})$ and then add $gs\mathcal{C}$ at the start and $\sigma(gs\mathcal{C})$ at the end and get a minimal gallery of chambers from $gs\mathcal{C}$ to $\sigma(gs\mathcal{C})$. By this argument and Theorem 2.19 any $\sigma$ stable apartment that contains $gs\mathcal{C}$ has to contain $g\mathcal{C}$. Therefore either case 1 or case 3 happen and we know that it isn’t case 3.
@@
Let $\mathcal{C}$ be the chamber fixed by $I$, let $g\in G$ be a representative of the orbit $x$ and let $\mathcal{A}$ be a $\sigma$ stable apartment that contains $g\mathcal{C}$ and $gs\mathcal{C}$.
In case number 1, by Lemma 5.5 all $\sigma$ stable apartments that contain $f=g\mathcal{C}\cap gs\mathcal{C}$ must contain $g\mathcal{C}$.
@@
This means that all chambers that contain $f$ and are different from $g\mathcal{C}$ are in the $H$ orbit of $gs\mathcal{C}$. The chamber $g\mathcal{C}$ itself is not in this $H$ orbit.
By Proposition 4.3 this yields $1_{x}(T_{s}+1)=1_{x\times s}+1_{x}$, or $1_{x}T_{s}=1_{x\times s}$.
@@
Let $x\in H\backslash G/I$ be as in the previous theorem, let $w\in W_{aff}(T)$ and let $l:W_{aff}\rightarrow\mathbb{N}$ be the usual length function on $W_{aff}$ (see Section 2.1 for a definition).
If $l_{\sigma}(x\times w)=2\cdot l(w)+l_{\sigma}(x)$ then
$1_{x}T_{w}=1_{x\times w}$
@@
Write $w=os_{1},...,s_{l(w)}$ as a product of simple reflections $s_{i}\in\tilde{\Delta}$ and $o\in\Omega$.
By Theorem 2.32 $T_{w}=T_{o}T_{s_{1}}...T_{s_{l(w)}}$.
For any $1\leq j\leq l(w)$ we have $l_{\sigma}(x\times o\times s_{1}\times....\times s_{j})\leq l_{\sigma}(x)+2j$ . We must have equality; otherwise, the equality in the assumption would fail, as acting by each $s_{i}$ can increase the length by at most $2$.
By Theorem 5.3, Proposition 4.1 and induction we obtain that $1_{x}T_{o}T_{s_{1}}...T_{s_{j}}=1_{x\times o\times s_{1}\times....\times s_{j}}.$ For $j=l(w)$, we obtain the required result. ∎
This result is enough to prove that $S(X)^{I}$ is finitely generated over ${\mathcal{H}}(G,I).$
@@
Let $T$ be a $\sigma$ stable maximal split torus of $G$ and let $\mathcal{A}$ be the apartment that corresponds to $T$. The torus $T$ acts on $\mathcal{A}$ by translations. Let $t\in T$ such that its action on $\mathcal{A}$ commutes with $\sigma$, then there is $t^{\prime}\in T\cap H$ whose action on $\mathcal{A}$ is the same as $t$.
@@
Let $\pi$ be a uniformizer of $\mathcal{O}$ and let $\Lambda(T)=\{\chi^{\vee}(\pi)|\chi^{\vee}\in X^{*}(T)\}$. The torus $T$ is a direct product of $\Lambda(T)$ and $T^{0}$ (see 3.6 in [12]). Write $t=t_{1}t_{2}$ with $t_{1}\in\Lambda(T)$ and $t_{2}\in T^{0}$.
Choosing a chamber $\mathcal{C}\subset\mathcal{A}$, we have $\sigma(t)\sigma(\mathcal{C})=\sigma(t\mathcal{C})=t\sigma(\mathcal{C})$, so $t^{-1}\sigma(t)$ acts as a translation that fixes $\sigma(\mathcal{C})$ and thus it acts trivially, hence $t^{-1}\sigma(t)\in T^{0}$. Thus $t_{1}^{-1}\sigma(t_{1})\in T^{0}$. The set $\Lambda(T)$ is $\sigma$ stable so we obtain $t^{-1}_{1}\sigma(t_{1})\in T^{0}\cap\Lambda(T)=1$ or $t_{1}\in H$. The elements $t$ and $t_{1}$ act the same way on $\mathcal{A}$ and the claim follows. ∎
@@
Let $\mathcal{A}$ be a $\sigma$ stable apartment and let $d\in\mathbb{N}$. The number of $H$ orbits of chambers $\mathcal{C}$ in $\mathcal{A}$ with $l_{\sigma}(\mathcal{C})=d$ is finite.